a ball thrown from the top of 10m high building with a horizontal speed of 5m/s. (a) calculate the time it takes to drop to the floor. (b) its horizontal displacement at the point when it reaches the floor


1
Expert's answer
2021-02-05T02:16:35-0500

(a)

"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot10\ m}{9.8\ \dfrac{m}{s^2}}}=1.43\ s"

(b)

"x=v_{0x}t=5\ \dfrac{m}{s}\cdot1.43\ s=7.15\ m."